The federal minimum wage has remained frozen for the past three years at​ $7.25 an​ hour, while the prices of gas and milk have risen steadily. Over this​ three-year period, the real value of the minimum wage has fallen to​ $6.77 per hour. By what percentage did the CPI increase over these three​ years?
Answer #1

Over this three-year period, the real value of the minimum wage has fallen to​ $6.77 per hour.
This means that at the end of the 3-year period, the real value of the minimum wage is
(6.77/7.25)*100 = 93.4%

So, decrease in real minimum wage = 100 - 93.4 = 6.6% (as in base year, real value of the minimum wage is 100%)
So, the change in real minimum wage 6.6%

Thus, increase in CPI must be 6.6%